OpenStudy (anonymous):

Two times the difference of a number and 6 is the same as five times difference of one less than a number. Find the number.

OpenStudy (anonymous):

2*(x-6)=5*(1-x)

OpenStudy (anonymous):

Simplifying 2x-12=5-5x

OpenStudy (anonymous):

7x=17

OpenStudy (anonymous):

x=17/7

OpenStudy (anonymous):

hi nano

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

hi iraqichick!! How r u??

OpenStudy (anonymous):

lol :D good u should go to the chat and introduce urself and meet some pplz :D

OpenStudy (anonymous):

lool yea after i finish my practice test lool

OpenStudy (anonymous):

good luck babez

OpenStudy (anonymous):

thank you hunni!!